# Ban&Pick: Achieving Free Performance Gains and Inference Speedup via Smarter Routing in MoE-LLMs

### v1

@@ 6.1 Experiment Setup
We conduct experiments on four fine-grained MoE models from two different model families, covering a wide range of parameter scales: DeepSeek-V2-Lite-Chat (16B), DeepSeek-V2.5-1210 (236B), Qwen3-30B-A3B, and Qwen3-235B-A22B. Our evaluation spans three categories of tasks using five commonly adopted datasets: AIME2024 (patel2024aimeaioptimizationmultiple, ) and Math-500 for mathematical reasoning, GPQA-Diamond (rein2023gpqagraduatelevelgoogleproofqa, ) for general knowledge reasoning, and LiveCodeBench (jain2024livecodebench, ) together with HumanEval+ (evalplus, ) for code generation. For DeepSeek-V2-Lite-Chat, we replace AIME2024 with GSM8K, as AIME2024 is overly difficult for this smaller model. We use OpenCompass (2023opencompass, ) as evaluation toolkit. The detailed generation hyperparameters are provided in the Appendix D.
@@ 7 Interaction Among Key Experts Across Tasks
From Table 4, we observe that enhancing multiple domains simultaneously still brings clear improvements compared to the baseline. For instance, when both Math and General experts are enhanced, AIME2024 increases by +4.0% and GPQA by +2.02%. However, these gains are slightly smaller than enhancing each domain individually ($0.66\%$ and $1.51\%$, respectively). When all three domains are enhanced together, the model continues to outperform the baseline—+3.33% on AIME2024, +2.02% on GPQA, and +3.05% on HumanEval+—though the gains are somewhat reduced compared to the best single-domain setting ($1.33\%$, $1.51\%$, and $1.22\%$). These results indicate that while domain-specific key experts do interfere to some extent, their combined enhancement remains clearly beneficial, confirming both their robustness and their crucial role in MoE models.
@@
This study is the first to reveal that in fine-grained MoE models, a very small set of domain‑specialized experts has a disproportionately large impact on final logits and task performance. By selectively increasing the utilization of these key experts and dynamically pruning redundant ones, we provide a novel and effective framework that improves both accuracy and efficiency. For future work, we suggest exploring automatic key expert selection without prior knowledge of the task type, as well as investigating more sophisticated enhancement strategies beyond uniform boosting.

### v2

@@ 6.1 Experiment Setup
We conduct experiments on four fine-grained MoE models from two different model families, covering a wide range of parameter scales: DeepSeek-V2-Lite-Chat (16B), DeepSeek-V2.5-1210 (236B), Qwen3-30B-A3B, and Qwen3-235B-A22B. Our evaluation spans three categories of tasks using five commonly adopted datasets: AIME2024 (patel2024aimeaioptimizationmultiple, ) and Math-500 for mathematical reasoning, GPQA-Diamond (rein2023gpqagraduatelevelgoogleproofqa, ) for general knowledge reasoning, and LiveCodeBench (V1) (jain2024livecodebench, ) with HumanEval+ (evalplus, ) for code generation. For DeepSeek-V2-Lite-Chat, we replace AIME2024 with GSM8K, as AIME2024 is overly difficult for this smaller model. We use OpenCompass (2023opencompass, ) as evaluation toolkit. The detailed generation hyperparameters are provided in the Appendix D.
@@ 7 Interaction Among Key Experts Across Tasks
From Table 4, we observe that enhancing multiple domains simultaneously still brings clear improvements. For instance, when both Math and General experts are enhanced, AIME2024 increases by +4.0% and GPQA by +2.02%. However, these gains are slightly smaller than enhancing each domain individually ($0.66\%$ and $1.51\%$, respectively). When all three domains are enhanced together, the model continues to outperform the baseline—+3.33% on AIME2024, +2.02% on GPQA, and +3.05% on HumanEval+—though the gains are somewhat reduced compared to the best single-domain setting. These results indicate that while domain-specific key experts do interfere to some extent, their combined enhancement remains clearly beneficial, confirming their crucial role.
@@
In this study, we reveal that in fine-grained MoE models, relatively balanced routing strategy underutilizes a small set of key experts which have a disproportionately large impact on task performance, and fixed expert activation introduces substantial redundancy. By selectively increasing the utilization of key experts and dynamically pruning redundant ones, we release the potential of routing and provide a novel and effective framework that improves both accuracy and efficiency. For future work, we suggest exploring the underlying mechanisms of how key experts enhance model performance, and investigating more sophisticated enhancement strategies beyond uniform boosting.